1)an object moves in a circular path of radius 10 m with a constant speed of 30m/s. What is the magnitude and the direction of the acceleration?

2.) a horizontal, 10m plank weighs 100N. it rests on two supports that are placed 1.0m from each end. how close to one end can an 800N person stand without causing the plank to tip?(hint: what happens to the normal force on the support at the opposite end of the plank from the man when the board tips?)
Answer #1

a) acceleration = V^2 /r

= (30)^2 /10 = 90 m/s2